# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 7. Probabilistic Approaches to the Energy Equality
Assume further that $\beta=\alpha/2.$ Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\alpha/2}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the kinetic energy equality holds:
@@
Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the $H^{\alpha-2\beta}$-energy equality holds:
@@
The probabilistic method breaks down when $\beta\neq\alpha/2.$ Instead, we will provide an analytic proof for this more general case, following a similar approach to the one used in [GHW24, Theorem 2.7].
@@
For any $\varepsilon>0$ and $m\in\mathbb{N}_{+}$, suppose that $\theta_{\varepsilon,m}$ is a solution of (3.7) with $\beta=\alpha/2$ and $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Then the law of $\theta_{\varepsilon,m}$ is invariant under $\mathfrak{T}$:
$\mathfrak{T}_{T}\theta_{\varepsilon,m}:=\left(-\theta_{\varepsilon,m}(T-t):0\leqslant t\leqslant T\right)\overset{\,\mathrm{d}}{=}\left(\theta_{\varepsilon,m}(t):0\leqslant t\leqslant T\right).$
@@
Let $\{e_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ be the Fourier basis of $H$ introduced in Section 2, and we set $M=M(m):=\sum_{0<|k|\leqslant m}1.$ The solution $\theta_{\varepsilon,m}$ can be expanded as $\theta_{\varepsilon,m}=\sum_{0<|k|\leqslant m}X_{k}e_{k}$, leading to the expression $u_{\theta_{\varepsilon,m}}=R^{\perp}\theta_{\varepsilon,m}=\sum_{0<|k|\leqslant m}X_{k}R^{\perp}e_{k}.$ We denote by $u_{k}:=R^{\perp}e_{k}$ and $B_{k,l,n}:=\langle u_{n}\cdot\nabla e_{l},e_{k}\rangle$, then (3.7) is equivalent to the following stochastic differential equations on $\mathbb{R}^{M}:$
[…]
Let $\mathcal{G}$ denote the Gaussian measure on $\mathbb{R}^{M}$ with mean zero and covariance $\varepsilon I/2$, where $I$ is the identity operator. Let $\mathcal{L}$ be the infinitesimal generator of the Markov process $X=(X_{k})_{0<|k|\leqslant m}$, decomposed as $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$, where
$\mathcal{L}_{1}f(x)$ $=\sum_{0<|k|\leqslant m}\left(-|k|^{2\alpha}x_{k}\partial_{k}f(x)+\frac{\varepsilon}{2}|k|^{2\alpha}\partial_{k}^{2}f(x)\right),$
$\mathcal{L}_{2}f(x)$ $=-\sum_{0<|k|\leqslant m}\sum_{0<|l|\leqslant m}\sum_{0<|n|\leqslant m}B_{k,l,n}x_{l}x_{n}\partial_{k}f(x),\quad\text{for all }f\in C_{b}^{2}\left(\mathbb{R}^{M}\right).$
For any $f,g\in C_{b}^{2}\left(\mathbb{R}^{M}\right)$, it follows from the integration by parts formula that
$\left\langle f,\mathcal{L}_{1}g\right\rangle_{L^{2}(\mathcal{G})}=\left\langle\mathcal{L}_{1}f,g\right\rangle_{L^{2}(\mathcal{G})},\quad\langle f,\mathcal{L}_{2}g\rangle_{L^{2}(\mathcal{G})}=-\langle\mathcal{L}_{2}f,g\rangle_{L^{2}(\mathcal{G})},$
which implies that $\mathcal{G}$ is an invariant measure of (7.3). The infinitesimal generator of $-X(\cdot)$ is $\mathcal{L}_{1}-\mathcal{L}_{2}$ and the infinitesimal generator of $X(T-\cdot)$ is $\mathcal{L}^{*}=\mathcal{L}_{1}^{*}+\mathcal{L}_{2}^{*}=\mathcal{L}_{1}-\mathcal{L}_{2}$, where denotes the adjoint operator in $L^{2}(\mathcal{G})$. This demonstrates the time-reversibility $\mathfrak{T}_{T}\theta_{\varepsilon,m}\overset{\mathrm{d}}{=}\theta_{\varepsilon,m}$.
@@
We begin by proving that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta).$ Suppose that $\varepsilon,m$ and $\theta_{\varepsilon,m}$ satisfy the conditions in Lemma 7.2. Let $\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\alpha/2}$ and let $\tilde{\mu}_{\varepsilon,m}:=\mu_{\varepsilon,m}\circ\mathfrak{T}_{T}^{-1}$. For any $\eta>0,$ it follows from Lemma 4.4 that there exists an open set $G$ in $\mathbb{X}_{\alpha,\alpha/2}$ such that $\theta\in G$ and $\mathcal{I}(\hat{\theta})>\mathcal{I}(\mathfrak{T}_{T}\theta)-\eta$ for all $\hat{\theta}\in\mathfrak{T}_{T}\overline{G}.$ Moreover, the continuity of $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ implies that $\mathfrak{T}_{T}^{-1}\overline{G}=\mathfrak{T}_{T}\overline{G}$ is closed in $\mathbb{X}_{\alpha,\alpha/2}.$ According to Proposition 6.3, Lemma 7.2 and Proposition 5.1, we have
$-\mathcal{I}(\theta)\leqslant\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)$ $=\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(G)$
$\leqslant\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(\overline{G})$
$=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(\mathfrak{T}_{T}\overline{G})$
$\leqslant-\inf_{\hat{\theta}\in\mathfrak{T}_{T}\overline{G}}\mathcal{I}(\hat{\theta})\leqslant-\mathcal{I}(\mathfrak{T}_{T}\theta)+\eta.$
@@
Corollary 7.3.
Let $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfy the scaling regime (2.8) and let $\theta_{\varepsilon,m},\mu_{\varepsilon,m}$ be as in Proposition 6.3. Suppose that $\theta\in\mathbb{X}_{\alpha,\alpha/2}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1.
(i) Assume that
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}>\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\theta).$
(ii) Assume that
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}<\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\mathfrak{T}_{T}\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\mathfrak{T}_{T}\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\mathfrak{T}_{T}\theta).$
Proof.
We first prove (i). Assume the converse, specifically that
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)\geqslant-\mathcal{I}(\theta).$ (7.6)
We will show that $\theta$ satisfies the kinetic energy inequality
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\leqslant\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$ (7.7)
and thereby obtain a contradiction. Let $\tilde{\mu}_{\varepsilon,m}:=\mu_{\varepsilon,m}\circ\mathfrak{T}_{T}^{-1}$. Using (7.6), Lemma 7.2 and Proposition 5.1, for any open set $G\subset\mathbb{X}_{\alpha,\alpha/2}$,
$-\mathcal{I}(\theta)\leqslant\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)$ $=\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(G)$
$\leqslant\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(\overline{G})$
$=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(\mathfrak{T}_{T}\overline{G})$
$\leqslant-\inf_{\hat{\theta}\in\overline{G}}\mathcal{I}(\mathfrak{T}_{T}\hat{\theta}).$
Since $\mathcal{I}$ is lower semi-continuous, $\sup_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\inf_{\hat{\theta}\in\overline{G}}\mathcal{I}(\mathfrak{T}_{T}\hat{\theta})=\mathcal{I}(\mathfrak{T}_{T}\theta).$ Therefore, $\mathcal{I}(\theta)\geqslant\mathcal{I}(\mathfrak{T}_{T}\theta)$. Then it follows from (7.4) and (7.5) that (7.7) holds. We can apply this argument again, with $\theta$ replaced by $\mathfrak{T}_{T}\theta$, to obtain (ii). ∎
@@
$\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\mathfrak{T}_{T}\theta,\tilde{g}\rangle\,\mathrm{d}s$ (7.9)
@@ 8.1. Conditional equivalence
$\langle\theta(t),\varphi(t)\rangle$ $=\langle\theta(0),\varphi(0)\rangle+\int^{t}_{0}\langle\theta,\partial_{s}\varphi\rangle\,\mathrm{d}s-\int^{t}_{0}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}s$ (8.2)
With the above definition, we can similarly define the weak-strong uniqueness class on $[0,\infty)$ as in Definition 6.1. Let
@@
Definition 8.2 (Weak-Strong Uniqueness Class on $[0,\infty)$).
We say $\mathcal{C}_{0}\subset\mathbb{X}^{\infty}_{\alpha,\beta}$ is a weak-strong uniqueness class of (8.1) for every $g\in L^{2}_{loc}([0,\infty);H)$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ of the skeleton equation (8.1) in the sense of Definition 8.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}^{\infty}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t>0,$
@@
Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $[0,\infty).$ Denote
$\mathcal{A}_{0}:=\big{\{}\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
[…]
For any $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ that satisfying $-\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, we define
@@
(i). It is sufficient to discuss the case that $\mathcal{U}(\phi)<\infty.$ In this case, there exists a function $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ such that $-\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, $\tilde{\theta}(0)=\phi$ and $\langle R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s),\tilde{\theta}(s)\rangle_{H^{\alpha-2\beta}(\mathbb{T}^{2})}=0$ for all $s<0.$ Then it follows from the definition of $\mathcal{A}_{0}$ that
@@
$S_{-\infty}(\tilde{\theta})$ $=\frac{1}{2}\int_{-\infty}^{0}\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}\,\mathrm{d}s+\int_{-\infty}^{0}\frac{\mathrm{d}}{\mathrm{d}s}\|\tilde{\theta}(s)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
@@
(ii). Now we aim to prove that the equality $\mathcal{U}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ holds as long as (8.5) admits a weak solution in $\mathcal{A}_{0}.$ Assume that $\bar{\theta}$ is the solution. Then
$\tilde{\theta}(\cdot):=-\bar{\theta}(-\cdot)$
satisfies $\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s)=0$ for all $s\leqslant 0$ and hence the first term on the right-hand side of (8.1) is 00. This completes the proof. ∎

### v3

@@ 7. Probabilistic Approaches to the Energy Equality
Assume further that $\beta=\alpha/2.$ Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\alpha/2}$ be a weak-strong uniqueness class of (1.1) such that $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ contains non-empty time-reversible subsets. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the kinetic energy equality holds:
@@
Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) such that $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ contains non-empty time-reversible subsets. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the $H^{\alpha-2\beta}$-energy equality holds:
@@
The probabilistic method breaks down when $\beta\neq\alpha/2.$ Instead, we will provide an analytic proof for this more general case, following a similar approach to the one used in [GHW24_LLNS, Theorem 2.7].
@@
For any $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, suppose that $\theta_{\varepsilon,m}$ is the solution of (3.1) with $\beta=\alpha/2$ and $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$. Then the law of $\theta_{\varepsilon,m(\varepsilon)}$ is invariant under $\mathfrak{T}$:
$\mathfrak{T}_{T}\theta_{\varepsilon,m(\varepsilon)}:=\left(-\theta_{\varepsilon,m(\varepsilon)}(T-t):0\leqslant t\leqslant T\right)\overset{\,\mathrm{d}}{=}\left(\theta_{\varepsilon,m(\varepsilon)}(t):0\leqslant t\leqslant T\right).$
@@
The proof is elementary. See [Tot20, Proposition 4.1] for details. ∎
@@
We begin by proving that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta).$ Suppose that $\varepsilon,m(\varepsilon)$ and $\theta_{\varepsilon,m(\varepsilon)}$ satisfy the conditions in Lemma 7.2. Let $\mu_{\varepsilon,m(\varepsilon)}$ be the law of $\theta_{\varepsilon,m(\varepsilon)}$ on $\mathbb{X}_{\alpha,\alpha/2}$ and let $\tilde{\mu}_{\varepsilon,m(\varepsilon)}:=\mu_{\varepsilon,m(\varepsilon)}\circ\mathfrak{T}_{T}^{-1}$. For any $\eta>0,$ it follows from Lemma 4.4 that there exists an open set $G$ in $\mathbb{X}_{\alpha,\alpha/2}$ such that $\theta\in G$ and $\mathcal{I}(\hat{\theta})>\mathcal{I}(\mathfrak{T}_{T}\theta)-\eta$ for all $\hat{\theta}\in\mathfrak{T}_{T}\overline{G}.$ Moreover, the continuity of $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ implies that $\mathfrak{T}_{T}^{-1}\overline{G}=\mathfrak{T}_{T}\overline{G}$ is closed in $\mathbb{X}_{\alpha,\alpha/2}.$ According to Proposition 6.1, Lemma 7.2, and Proposition 5.1, we have
$-\mathcal{I}(\theta)\leqslant\liminf_{\varepsilon\rightarrow 0}\varepsilon$ $\log\mu_{\varepsilon,m(\varepsilon)}(G)=\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m(\varepsilon)}(G)\leqslant\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m(\varepsilon)}(\overline{G})$
$=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m(\varepsilon)}(\mathfrak{T}_{T}\overline{G})\leqslant-\inf_{\hat{\theta}\in\mathfrak{T}_{T}\overline{G}}\mathcal{I}(\hat{\theta})\leqslant-\mathcal{I}(\mathfrak{T}_{T}\theta)+\eta.$
@@
Proof of Corollary 1.5.
The proof follows from an argument analogous to [GHW24_LLNS, Remark 2.8], which is similar in spirit to the probabilistic proof of Theorem 7.1. Thus we omit it. ∎
@@
$\quad\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\mathfrak{T}_{T}\theta,\tilde{g}\rangle\,\mathrm{d}s$ (7.6)
@@ 8.1. Conditional equivalence
$\langle\theta(t),\varphi(t)\rangle$ $=\langle\theta(0),\varphi(0)\rangle-\int^{t}_{0}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}s-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s$ (8.2)
$+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s+\int^{t}_{0}\langle\theta,\partial_{s}\varphi\rangle\,\mathrm{d}s+\int^{t}_{0}\langle\Lambda^{2\beta}\varphi,g\rangle\,\mathrm{d}s.$
With the above definition, we can similarly define the weak-strong uniqueness class on $0,\infty)$ as in Definition 1.1. Let
@@
Definition 8.2 (Weak-Strong Uniqueness Class on $0,\infty)$).
We say that $\mathcal{C}_{0}\subset\mathbb{X}^{\infty}_{\alpha,\beta}$ is a weak-strong uniqueness class of (8.1) if for every control $g\in L^{2}_{loc}([0,\infty);H)$, the following holds: for arbitrary two weak solutions $\theta_{1},\theta_{2}$ of the skeleton equation (8.1) in the sense of Definition 8.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}^{\infty}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t>0$,
@@
Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $0,\infty).$ Denote
$\mathcal{A}_{0}=\big\{\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
$\lim\limits_{t\rightarrow\infty}\|\theta(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}=0$ $\big\}.$
For any $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ that satisfying $\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, we define
@@
(i). It is sufficient to discuss the case that $\mathcal{U}(\phi)<\infty.$ In this case, there exists a function $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ such that $\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, $\tilde{\theta}(0)=\phi$ and $\langle R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s),\tilde{\theta}(s)\rangle_{H^{\alpha-2\beta}(\mathbb{T}^{2})}=0$ for all $s<0.$ Then it follows from the definition of $\mathcal{A}_{0}$ that
@@
[…]
$=\frac{1}{2}\int_{-\infty}^{0}\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}\,\mathrm{d}s+\int_{-\infty}^{0}\frac{\mathrm{d}}{\mathrm{d}s}\|\tilde{\theta}(s)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
@@
(ii). Now we aim to prove that the equality $\mathcal{U}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ holds as long as (8.5) admits a weak solution in $\mathcal{A}_{0}.$ Assume that $\bar{\theta}$ is the solution. Then $\tilde{\theta}(\cdot):=-\bar{\theta}(-\cdot)$ satisfies $\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s)=0$ for all $s\leqslant 0$ and hence the first term on the right-hand side of (8.1) is 0. This completes the proof. ∎